# A Decision Theoretic Framework for Measuring AI Reliance

### v4

@@ 5.3.1. Experiment Design
The experiment compares AI-assisted human recommendations under two different conditions: anchoring and non-anchoring. Participants assigned to the anchoring condition see the question presented together with the AI’s recommendation, while under the non-anchoring condition, participants are asked to predict the risk before seeing AI recommendation and then to revise their assessment after having AI recommendation. In each question, participants are shown the profile of a defendant, including demographics, current charge, and criminal history. Participants are asked to report: 1) the probability of the defendant being re-arrested from $[0,100\%]$, and 2) a binary choice of whether the defendant will be re-arrested within a given duration or not. The decision tasks for probabilty and binary decision are summarised in Table 5 in Appendix B.
@@ 6. Discussion
We contribute a formal definition of reliance and corresponding framework for interpreting losses in behavioral decision-making performance within the baseline and benchmark for complementary performance. The first source of loss concerns the difference in the rate at which the behavioral decision-maker relies on the AI relative to the appropriate level of reliance defined by the decision problem, calculated in payoff space. The second source of loss concerns the difference in score between a behavioral decision-maker and the best score a rational decision-maker who relies on the AI at the same rate as the behavioral decision-maker but who perfectly perceives the posterior probabilities could achieve. By contributing clear comparison points in the form of performance benchmarks to the design and interpretation of studies of human reliance on AI, our work enables researchers to identify the upper-bound of complementary performance and how far the human-AI team is from this optimal attainable performance.
Closest to the motivation of our work, Fok and Weld (Fok and Weld, 2023) motivate the need for a notion of “strategy-graded reliance,” where appropriate reliance is determined from the relative expected performance of the human and the AI, over “outcome-graded reliance” based on the human’s acceptance of AI advice conditioned on its post-hoc correctness. Several other prior works propose studying reliance using conditional probability (e.g., (Schemmer et al., 2023; Yang et al., 2020; Wang and Yin, 2021; Schoeffer et al., 2023; Wang et al., 2022; Ma et al., 2023)) to separate cases where the human recommendation is better than the AI recommendation from cases where the AI recommendation is better than the human recommendation. We unambigously define strategy-guided reliance and show how to calculate optimal reliance and disentangle sources of behavioral loss.
@@ Appendix A The algorithms for calculations in the framework
for $action\leftarrow\{0,1\}$ ($action=0$ follow human, $1$ follow AI) do
@@
Input: the experimental data $D$ with each row representing one experimental trial, the space of derived binary-adoption states $\hat{\Theta}$, and the space of signals $V$
Output: the empirical distribution $\tilde{\pi}(\hat{\theta},v)$
@@
/* Initializing a matrix with all 00. */
@@
       $v_{i}\leftarrow\text{the signal realized in }row_{i}$;
       $\tilde{\pi}(\hat{\theta}_{i},v_{i})\leftarrow\tilde{\pi}(\hat{\theta}_{i},v_{i})+1$;
@@
$\tilde{\pi}(\hat{\theta},v)\leftarrow\tilde{\pi}(\hat{\theta},v)/|\tilde{\pi}(\hat{\theta},v)|$;
@@
Output: the empirical distribution $\tilde{\pi}(\hat{\theta},\tilde{v})$ on the optimally discretized space
@@
/* Initializing a matrix with all 00. */
$\{v_{i}\}\leftarrow\text{all signals realized in }D$;
$kmeans\leftarrow initialize\_kmeans(\{v_{i}\},K)$;
@@
       $v_{i}\leftarrow\text{the signal realized in }row_{i}$;
       $\tilde{v}_{i}\leftarrow kmeans(v_{i})$;
@@
Input: the experimental data $D$ with each row representing one experimental trial, the joint distribution between states and signals $\pi(\hat{\theta},v)$, and the scoring rule for the derived binary-adoption decision task $\hat{S}$
@@
       $v\leftarrow\text{the signal realized in }row$;
       $\pi(\hat{\theta}|v)=\pi(\hat{\theta},v)/\pi(v)$;
@@
       $action\leftarrow\operatorname*{argmax}_{\hat{a}\sim\{\text{human},\text{AI}\}}{E_{\hat{\theta}\sim\pi(\hat{\theta}|v)}(\hat{S}(\hat{a},\theta))}$;
@@
Input: the experimental data $D$ with each row representing one experimental trial, the joint distribution between states and signals $\pi(\hat{\theta},v)$, the scoring rule for the original decision task $S$, and the scoring rule for the derived binary-adoption decision task $\hat{S}$
@@
       Sort $\text{Sort }P_{i}\text{ in decreasing order of }\text{\bf E}_{\pi(\hat{\theta}|v)}\![{S(y^{AI},y)}]-\text{\bf E}_{\pi(\hat{\theta}|v)}\![{S(y^{H},y)}]$;
       $\{v_{j}\}\leftarrow\{\text{the signal realized in }row_{j}\}_{row_{j}\in P_{i}}$;
@@
       $\{a^{b}_{j}\}\leftarrow\{\text{action drawn from }\pi(a^{b}|\theta_{j},v_{j})\}_{row_{j}\in P_{i}}$;
@@
Input: the experimental data $D$ with each row representing one experimental trial, total number of iterations $T$, the sample size $k$, prior distribution of the binary-adoption state $\pi(\hat{\theta})$, the joint distribution between states and signals $\pi(\hat{\theta},v)$, the joint behavior $\pi(v,\theta,a^{b})$, the scoring rule $S$, and the scoring rule for derived binary-adoption decision task $\hat{S}$
@@
       $\textsc{R}_{\varnothing}{}_{i}\leftarrow\text{Rational baseline}(\tilde{D},\pi(\hat{\theta}),\hat{S})$;
       $\textsc{R}_{i}\leftarrow\text{Rational benchmark}(\tilde{D},\pi(\hat{\theta},v),\hat{S})$;
       $\textsc{B}_{i}\leftarrow\text{Behavioral performance}(\tilde{D},\pi(v,\hat{\theta},a^{b}),S)$;
       $\textsc{R}^{\text{m}}_{i}\leftarrow\text{Mis-reliant rational baseline}(\tilde{D},\pi(\hat{\theta},v),S,\hat{S})$;
@@ Appendix B Formalized decision tasks
Action (choice) $a\in\{0,1\}$ Be re-arrested or not $\hat{a}\in\{0=\text{human},1=\text{AI}\}$
Signal $v=\{x,y^{H},y^{AI},e(y^{AI})\}$
@@
Action (choice) $a\in\{0,1\}$ Deceptive or genuine $\hat{a}\in\{0=\text{human},1=\text{AI}\}$
Signal $v=\{x,y^{H}(x),y^{AI}(x),e(y^{AI})\}$
@@
Action (choice) $a\in\{0,1\}$ Re-arrest or not $a\in[0,1]$ Probability of re-arrest $\hat{a}\in\{0=\text{human},1=\text{AI}\}$
Signal $v=\{x,y^{H}(x),y^{AI}(x),e(y^{AI})\}$

### v5

@@ 5.3.1. Experiment Design
The experiment compares AI-assisted human recommendations under two different conditions: anchoring and non-anchoring. Participants assigned to the anchoring condition see the question presented together with the AI’s recommendation, while under the non-anchoring condition, participants are asked to predict the risk before seeing AI recommendation and then to revise their assessment after having AI recommendation. In each question, participants are shown the profile of a defendant, including demographics, current charge, and criminal history. Participants are asked to report: 1) the probability of the defendant being re-arrested from $0,100\%]$, and 2) a binary choice of whether the defendant will be re-arrested within a given duration or not. The decision tasks for probabilty and binary decision are summarised in Table 5 in Appendix B.
@@ 6. Discussion
We contribute a formal definition of reliance and corresponding framework for interpreting losses in behavioral decision-making performance within the baseline and benchmark for complementary performance. The first source of loss concerns the difference in the rate at which the behavioral decision-maker relies on the AI relative to the appropriate level of reliance defined by the decision problem, calculated in payoff space. The second source of loss concerns the difference in score between a behavioral decision-maker and the best score a rational decision-maker who relies on the AI at the same rate as the behavioral decision-maker but who perfectly perceives the posterior probabilities could achieve. By contributing clear comparison points in the form of performance benchmarks to the design and interpretation of studies of human reliance on AI, our work enables researchers to identify the upper-bound of complementary performance and how far the human-AI team is from this optimal attainable performance.context
Closest to the motivation of our work, Fok and Weld (Fok and Weld, 2023) motivate the need for a notion of “strategy-graded reliance,” where appropriate reliance is determined from the relative expected performance of the human and the AI, over “outcome-graded reliance” based on the human’s acceptance of AI advice conditioned on its post-hoc correctness. Several other empirical works propose studying reliance using conditional probability (e.g., (Schemmer et al., 2023; Yang et al., 2020; Wang and Yin, 2021; Schoeffer et al., 2023)) to separate cases where the human recommendation is better than the AI recommendation from cases where the AI recommendation is better than the human recommendation. Information display in human-AI collaboration also has extended to using information of human accuracy before receiving AI (Wang et al., 2022; Ma et al., 2023), suggesting studies about complementarity to consider correctness likelihood of both human and AI. We unambigously define strategy-guided reliance and show how to calculate optimal reliance and disentangle sources of behavioral loss.
@@ Appendix A The algorithms for calculations in the framework
for $action\leftarrow\{0,1\}$ ($action=0$ follow human, 11 follow AI) do
@@
Input: the experimental data $D$ with each row representing one experimental trial, the space of derived binary-adoption states $\hat{\Theta}$, and the space of signals $\hat{V}$
@@
$\tilde{\pi}(\hat{\theta},\hat{v})\leftarrow 0\,{\mathds{1}}_{|\hat{\Theta}|}{\mathds{1}}_{|\hat{V}|}^{\top}$;
/* Initializing a matrix with all 0. */
@@
 $\hat{v}_{i}\leftarrow\text{the signal realized in }row_{i}$;
 $\tilde{\pi}(\hat{\theta}_{i},\hat{v}_{i})\leftarrow\tilde{\pi}(\hat{\theta}_{i},\hat{v}_{i})+1$;
@@
[…]
@@
$\tilde{\pi}(\hat{\theta},\tilde{v})\leftarrow 0\,{\mathds{1}}_{|\hat{\Theta}|}{\mathds{1}}_{K}^{\top}$;
/* Initializing a matrix with all 0. */
$\{\hat{v}_{i}\}\leftarrow\text{all signals realized in }D$;
$kmeans\leftarrow initialize\_kmeans(\{\hat{v}_{i}\},K)$;
@@
 $\hat{v}_{i}\leftarrow\text{the signal realized in }row_{i}$;
 $\tilde{v}_{i}\leftarrow kmeans(\hat{v}_{i})$;
@@
[…]
@@
 $\hat{v}\leftarrow\text{the signal realized in }row$;
[…]
@@
[…]
@@
[…]
@@
 $\text{Sort }P_{i}\text{ in decreasing order of }\text{\bf E}_{\pi(\hat{\theta}|\hat{v})}\![{S(y^{AI},y)}]-\text{\bf E}_{\pi(\hat{\theta}|\hat{v})}\![{S(y^{H},y)}]$;
 $\{\hat{v}_{j}\}\leftarrow\{\text{the signal realized in }row_{j}\}_{row_{j}\in P_{i}}$;
@@
 $\{a^{b}_{j}\}\leftarrow\{\text{action drawn from }\pi(a^{b}|\theta_{j},\hat{v}_{j})\}_{row_{j}\in P_{i}}$;
@@
[…]
@@
[…]
 $\textsc{B}_{i}\leftarrow\text{Behavioral performance}(\tilde{D},\pi(v,\theta,a^{b}),S)$;
[…]
@@ Appendix B Formalized decision tasks
This appendix includes the formalized decision tasks for the experiments in the main text.
@@
[…]
Signal $v=\{x,y^{AI},e(y^{AI})\}$ $\hat{v}=\{x,y^{H},y^{AI}\}$
@@
[…]
Signal $v=\{x,y^{AI},e(y^{AI})\}$ $\hat{v}=\{x,y^{H},y^{AI}\}$
@@
Note that the feature $x$ is included in the binary-adoption signals in Table˜4 because $y^{H}$ and $y^{AI}$ in Lai and Tan (2019) are both binary-valued, which makes the signal space too small to generate a meaningful benchmark.
@@
[…]
Signal $v=\{x,y^{AI},e(y^{AI})\}$ $\hat{v}=\{x,y^{H},y^{AI}\}$